ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

Round Table

1991 MS Classification: 0 Kersten, Ina, Universität Bielefeld International Comparison of Mathematical Studies, University Degrees and Professional Perspectives August 25, 1998, 7.30 pm, H 104.

International cooperation and collaboration is increasing. However, the rules and regulations in many countries seem to impede the international mobility of students, teachers, and professionals. A discussion of this and related topics -with respect to mathematics – will be the subject of a round table organized by Ina Kersten (Bielefeld, Germany). Among the panelists are Doina Cioranescu (Laboratoire d’Analyse Numérique, Univ. Pierre et Marie Curie, Paris, France), Karl-Heinz Hoffmann (President of the German Mathematical Society, formerly the President of the “Wissenschaftsrat” of Germany), Jürgen Netzel (Labour Exchange, Berlin, Germany), Andrzej Pelczar (President of the General Council for Higher Education in Poland, Vice-President of the European Mathematical Society) and Mario Primicerio (Mayor of Florence, Italy). There will be short reports on higher education policies and professional perspectives in different countries. The discussion also includes mobility of students and teaching staff, and the development of cooperation networks. There are many professional fields where mathematicians are needed as the discussion will show.